OpenStudy (anonymous):

What values of x and y satisfy and simultaneously? x = 3 and y = −10 x = −3 and y = −3 x = 1 and y = 5 x = 3 and y = 10

OpenStudy (mrnood):

you need to post the graph or equations

OpenStudy (anonymous):

-\[-13/3x-2y=7 and 3x+2/5y=5\]

OpenStudy (anonymous):

thats the equation given

OpenStudy (mrnood):

you probably need some brackets for clarification - or use the equation editor. is it \[\frac{ -13 }{ 3 }x -2y =7\] \[3x+\frac{ 2 }{ 5 }y = 5\]

OpenStudy (anonymous):

yes im new to this so i didnt know how to put the brackets but thats correct

OpenStudy (mrnood):

do you know the elimination method for solving simultaneous equations? try multiplying the second equation by 5, and post the new equation here